OpenStudy (anonymous):

8 (20x11) x (12+48) divided by 6 - (9-2)

OpenStudy (anonymous):

the answer is 216 in the book it says.. but im not sure how to find this out. i keep getting 713 for my answer

OpenStudy (jdoe0001):

what's 8(20x11) x (12+48)?

OpenStudy (jdoe0001):

$$ \cfrac{8(20\times 11)\times (12+48)}{6-(9-2)} \huge \ ? $$

OpenStudy (anonymous):

yep thats the problem

OpenStudy (anonymous):

im just not sure how to solve it. i keep getting the wrong answer over and over again

OpenStudy (jdoe0001):

well, I sure don't get 713

OpenStudy (jdoe0001):

the bottom part will be 6-(7) = -1 so whatever you get atop, just negativize it

OpenStudy (jdoe0001):

ohh, you said the book says is 216, well, I don't get that either

OpenStudy (anonymous):

wait i typed it wrong

OpenStudy (anonymous):

8 x (20-11) x (12+48) divided by 6 - (9-2)

OpenStudy (anonymous):

that is the correct problem now. sorry about that

OpenStudy (jdoe0001):

all numbers atop are positive, the denominator is negative the answer will surely be a negative number, no 216 though the numerator can be narrowed down to 8 x 220 x 60 and that's a really big figure, much bigger than 216

OpenStudy (anonymous):

8x9x60 divied by 6 - 7 is 713

OpenStudy (anonymous):

that is the wrong problem i typed..

OpenStudy (jdoe0001):

ahemm, 6-7 is a denominator in itself, so that's -1, 6-7

OpenStudy (jdoe0001):

9? I gather you meant to use 20-11, as opposed to 20 x 11 then

OpenStudy (anonymous):

ok but its not 8x220

OpenStudy (anonymous):

it would be 8x9x60 so i dont know why the back of my book is telling me wrong answers. so frustrating

OpenStudy (jdoe0001):

can you take a screenshot of the actual material?

OpenStudy (anonymous):

it makes no sense why its saying the answer is 216 hmmm

OpenStudy (anonymous):

my webcam isnt that great i just tried

OpenStudy (jdoe0001):

well, something tells me what you typed isn't what the paper is suggesting

OpenStudy (anonymous):

im just going to go to the help center before class for this problem because the idiots who wrote this book dont know how to solve problems correctly apparently and then confused students